All i know is the question, 1+ the square root of (x+5) this is equal to the square root of (x+12) Now what is x??
Answer by checkley77(12844)

1+sqrt(x+5)=sqrt(x+12) square botgh sides.
1+2sqrt(x+5)+x+5=x+12
x-x+2sqrt(x+5)=12-5-1
2sqrt(x+5)=6 square both sides.
4(x+5)=36
4x+20=36
4x=36-20
4x=16
x=16/4
x=4 ans.
Proof:
1+sqrt(4+5)=sqrt(4+12)
1+sqrt9=sqrt16
1+3=4
4=4